Sally travels 67m south 45m west find the displacement and distance?
1
Expert's answer
2020-03-16T13:05:01-0400

The distance


"d=67+45=112\:\rm m."

The displacement

"s=\sqrt{67^2+45^2}=81\:\rm m."